OpenStudy (anonymous):

Describe the slope of the tangent line to the curve defined by x 2 + xy - y 2 = 5 when x = 2. positive negative zero undefined

OpenStudy (dumbcow):

find dy\dx and plug in x=2

OpenStudy (ranga):

do implicit differentiation and find y'. put x = 2.

OpenStudy (anonymous):

2x+ dy/dx (x*y)dy/dx 2y=0

OpenStudy (anonymous):

2x+dy/dx (x*y) -2y=0

OpenStudy (ranga):

I prefer using y' when doing implicit differentiation so it is less clutter than using dy/dx.

OpenStudy (anonymous):

did I do that right?

OpenStudy (ranga):

x^2 + xy - y^2 = 5 2x + xy' + y - 2yy' = 0 put x = 2 right away to get rid of the x terms. 4 + 2y' + y - 2yy' = 0 4 + y + y'(2 - 2y) = 0 y' = -(4 + y)/(2 - 2y) We now need to find y when x = 2. Put x = 2 in the original equation and solve for y. Then put those y values in y' and find the slope.

OpenStudy (anonymous):

4+2y-y^2=5?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

okay is that all I do?

OpenStudy (anonymous):

How do I know if it is positive or negative?

OpenStudy (ranga):

solve the quadratic for y first.

OpenStudy (anonymous):

how do I do that?

OpenStudy (ranga):

not familiar with solving quadratic equations?

OpenStudy (anonymous):

oh sorry I know how to do that. I am really tired and I am trying to study for a test

OpenStudy (anonymous):

I studied like 7 1/2 hours yesterday and I just have everything confused. Sorry about that.

OpenStudy (ranga):

You got 4+2y-y^2=5 4 + 2y - y^2 - 5 = 0 -1 + 2y - y^2 = 0 multiply by -1 y^2 - 2y + 1 = 0 factor (y - 1)^2 = 0 y = 1. So when x = 2, y = 1. Put y = 1 in y' y' = -(4 + y)/(2 - 2y) y' = -(4 + 1) / (2 - 2) = -5 / 0 (division by zero) So slope is UNDEFINED at x = 2. That means the tangent to the curve at x = 2 is a vertical line.

OpenStudy (anonymous):

Do you use the same rules for any problem like this?

OpenStudy (anonymous):

or steps.

OpenStudy (ranga):

If you are given a function or an equation and you are asked to find the slope at a point you will have to find y' or dy/dx and evaluate y' at the given x. But if you do implicit differentiation the y will all be mixed up in the y' equation and so you will have to find the y value at the given x which you can do by putting x in the original equation and solving for x. Then you put that y in y' and you will have your slope.

OpenStudy (anonymous):

Thanks so much for helping me!

OpenStudy (ranga):

you are very welcome.